# The search for Population III: Confirmation of a HeII emitter with no metal lines at z=10.6

### 20 Mar 2026 v1

@@ Abstract
The detection and characterization of the first population of stars, formed out of pristine clouds of gas, the so-called Population III (PopIII), is one of the most important goals of modern astrophysics. These stars are theorised to form in galactic embryos within the first few hundred million years after the Big Bang (e.g., see Abel et al., 2002; Yoshida et al., 2003; Klessen and Glover, 2023, for a recent review). There have been detections of galaxies with extremely low metallicities, in the range (e.g., Vanzella et al., 2023; Fujimoto et al., 2025; Hsiao et al., 2025; Maiolino et al., 2025; Morishita et al., 2025; Nakajima et al., 2025). These could be very low metallicity PopII galaxies, or possibly self-polluted or hybrid PopIII galaxies (Rusta et al., 2025). However, these metallicities are still well above the expected threshold for PopIII, which are generally identified below a critical metallicity , which prevents the formation of low-mass stars because of inefficient gas cooling (e.g., Bromm et al., 2002; Schneider et al., 2002; Bromm and Yoshida, 2011).
@@
In this paper, we report deep high-resolution (G235H; ) NIRSpec IFU observations of the same region targeted in the previous study. These new observations confirm the presence of HeII emission in the halo of GN-z11. Further confirmation is obtained through the detection of H at the same location and same redshift, which is reported in a companion paper (Übler et al. 2026).
A second companion paper, Rusta et al. (submitted), presents a theoretical interpretation of our findings on this source in the Pop III scenario, also discussing the implications for the still-unknown mass distribution of Pop III stars.
@@ 2 Observations and data processing
Observations were obtained through the JWST programme ID 5086 using the NIRSpec IFU (Jakobsen et al., 2022; Böker et al., 2022, 2023) with the high resolution grating G235H (R2700). We recall that IFU has a field of view of . The observations were performed between May 15 and 17, 2025, split into two visits. Detectors were read out with the NRSIRS2 pattern, with 25 groups and 2 integrations. A total of 39 dithers were executed, with a medium cycling pattern. Unfortunately, during dither position 18 the telescope guiding camera locked onto a hot pixel; hence, this frame had to be discarded in further processing. After removing this frame, the total exposure time was 38.8 hours. The data reduction uses the JWST Science calibration pipeline v1.15.0 with CRDS jwst_1293.pmap. To improve the quality of the datacube, we performed a number of further data processing steps, in addition to the default pipeline. The customised procedures used for residual cosmic rays snowballs flagging, subtraction of pink noise (1/f noise), and removal for failed open MSA shutters signal are described in detail in Perna et al. (2023). Stage 2 was executed using the bad-pixel self-calibration implemented in the standard pipeline, in which all exposures on a given detector are used to identify and flag bad pixels that may have been missed by the reference bad-pixel mask. For the final stage of the pipeline, outlier rejection was performed following D’Eugenio et al. (2024b), with a rejection threshold at the percentile. Before combination, the cube was resampled to a scale of per spaxel.
@@ 3 Hebe: confirmation of a HeII emitter in the halo of GN-z11
The upper panel of Fig.1 shows the spectrum extracted from a aperture centred at East and North of GN-z11 (RA 12:36:25.520, DEC +62:14:31.934). The spectrum shows an emission feature, detected at 6, which is identified as HeII1640 at (centroid of the line), blue-shifted by 450 km/s relative to GN-z11. Note that, for this paper, the redshift of GN-z11 is consistently remeasured based on its HeII emission in our cube, and found to be ; this is slightly lower than the redshift given by Bunker et al. (2023) using medium resolution spectroscopy - - but it is also slightly higher than the redshift that would be derived from their HeII medium resolution detection, which would give ; these slight inconsistencies are likely due to the lower spectral resolution of the previous observations.
This independently confirms the previous HeII detection at the same location obtained with medium resolution spectroscopy by Maiolino et al. (2024b). We name Hebe this HeII emitting source, which is located at 3 kpc from GN-z11.
@@
Fig.2 shows the HeII map by collapsing the continuum-subtracted spectral channels between 1.8995 m and 1.9007 m (the three central channels of the line), which highlights the location of the HeII emitter, which, as already mentioned, we name Hebe. The red star indicates the location of GN-z11. Fig.3 shows the HeII contours again, but overlaid on the NIRCam RGB image, with the Field of View adjusted to show the location of Hebe relative to GN-z11 and a foreground galaxy at z=2.04, which causes some gravitational lensing (see Appendix C).
@@
As discussed in more detail in Section 6 below, the location and flux of the HeII emission are broadly consistent with the previous detection based on the NIRSpec medium resolution grating (Maiolino et al., 2024b), although the comparison with the latter implies the presence of some diffuse emission, both spatially and in velocity. Fig.2 shows the compactness of the core of the line emission, which is essentially unresolved ( 400 pc in size) at the NIRSpec resolution. We note some diffuse emission extending towards the SW. Even though its surface brightness is not statistically significant (mostly within of the background fluctuations), a similar diffuse emission was seen in the medium resolution map. This will be discussed again in Section 6. There is also another marginally detected () emitter 0.8′′ North of GN-z11; this was also marginally detected in medium resolution map, and will be discussed in Section 6.
@@
The HeII emission has a FWHM of 110 km/s (deconvolved by the spectral resolution)222Following, e.g., D’Eugenio et al., 2025 or Shajib et al., 2025, we take the spectral resolution as the nominal value (Jakobsen et al., 2022), but at 1.9 m this is multiplied by a factor of 0.7 (determined empirically in those references), which takes into account that, for compact sources, the size of the source can be comparable to the size of the IFU slices., and is marginally resolved in two components. This is in contrast with the H emission, which is spectrally marginally unresolved (see Übler et al. 2026). We have simultaneously fitted HeII and H with two Gaussian components (C1 and C2), as shown by the dashed lines in Fig.1. The two components are separated by km/s. The simultaneous fit indicates that H is associated only with C2, while the component C1 of H is undetected.
@@
We have also attempted a simultaneous fit with a single Gaussian component, which results in a significantly worse fit, as shown in Appendix A, with .
@@
It is difficult to explore whether there is any small spatial offset between the two components, as the maps of individual spectral channels become too noisy for a proper assessment.
@@
The inferred fluxes and luminosities of the HeII and H lines are reported in Table 1. These include an aperture correction factor of 3.03, estimated from the cubes of two stars (ID 1219 –PI: Luetzgendorf – and ID 1222 –PI: Willott–) processed with the same methodology, at the wavelength of 1.9m. The total flux has been estimated by simply summing up the flux in the central 5 spectral pixels of the line. The flux of the individual components is determined by the simultaneous double Gaussian fit.
@@
As discussed in Section 6, there is evidence for a diffuse component of HeII; in the second part of Table 1 we also provide an estimate of the HeII’s emitter flux after subtracting this more diffuse component333This is done by taking the diffuse flux estimated in Sect.6 and deriving the contribution to the extraction aperture..
@@ 4 Equivalent width constraints
Measuring or constraining the equivalent width of the HeII emission is incredibly important, as it can provide key constraints on the nature of the source. This requires a detection of the continuum emission. For extremely low metallicity galaxies, the main difficulty is the continuum detection rather than the nebular line emission. A clear example is the extremely metal poor galaxy LAP-1 at (Vanzella et al., 2023), where multiple nebular emission lines are clearly detected, while the continuum remains undetected despite deep exposures and lensing magnification (Nakajima et al., 2025). The HeII emitter in GN-z11 has a similar issue, with the additional complication that it is in the proximity of the relatively extended foreground galaxy at z=2.028. As discussed in Maiolino et al. (2024b) (see also Tacchella et al., 2023), modelling and removing the emission from the foreground galaxy leaves some residual flux around the location of Hebe. Yet, there is no clear indication of a continuum source associated with it, suggesting that the flux is some diffuse residual after the foreground galaxy removal. Thanks to the more precise localization of Hebe, we can now confirm the absence of a continuum detection at the 5 upper limit level of 3.1 nJy in F200W. We derive this value from the same aperture adopted for the spectrum extraction, accounting for a correction for a point-like source. Direct photometry (also corrected for the limited aperture size) provides a flux measurement of nJy. The 5 upper limit was corrected for the contribution of the HeII line to the broad-band measurement to get a spectral continuum upper limit. The rest-frame EW(HeII) corresponding to this continuum is 47 Å. We provide a conservative lower limit on the EW of the C1 and C2 components by applying the same continuum upper limit to each of them. This translates into lower limits of 28 Å and 22 Å, for the C1 ad C2 components, respectively. All these measurements are provided in Table 1.
As we discuss extensively in Section 7, the inferred rest-frame EW(HeII) are all above 20Å, well above the EW expected for a metal-enriched (PopII) stellar population.
@@ 5 Absence of metal emission lines
The G235H/F170LP setup covers a wavelength range where various metal emission lines typically detected in high-redshift galaxies (e.g. D’Eugenio et al., 2024a; Scholtz et al., 2025a) are expected to be seen at the redshift of the HeII emitter. Specifically, the NIV]1483,1486, CIV1548,1551, OIII]1661,1666, and CIII]1906,1908 doublets, as well as the NIII]1749 multiplet, are all within the spectroscopic range of this setup. Considering that in local and high-z galaxies at least some these lines (if not all of them) are stronger than HeII, at least some of these metal lines should be visible also in Hebe. On the contrary, none of these transitions is detected in Hebe. We report conservative upper limits of the flux of these metal lines relative to HeII. The spectral regions around the location of these lines are shown in Appendix A.
@@
The companion paper by Rusta et al. (subm.) provides a theoretical interpretation of our observational findings, showing that these non-detections are very constraining of the nature of the ionizing sources.
@@ 6 Consistency with the medium resolution observation and diffuse HeII emission
Secondly, in Fig.4 we compare the HeII map obtained in this work, with the HeII map obtained in Maiolino et al. (2024b) (contours). Note that the G235M map shows significant HeII emission associated with GN-z11, not seen in the G235H map because the velocity range integrated in G235M is larger and includes also some emission from GN-z11 (recall that Hebe is blueshifted by km/s relative to GN-z11). We have astrometrically registered the two images via the GN-z11 continuum. Unfortunately, this leaves some uncertainty in the relative positions. Specifically, in addition to the coarseness of the NIRSpec spatial sampling ( native spaxels), there are only two continuum sources that can be used for the alignment: GN-z11 and the foreground galaxy, which is not point-like but fairly extended. The white circle in the figure illustrates the location of GN-z11, and its size illustrates the uncertainty in aligning the two continua. The HeII emission in G235H appears slightly offset towards the NE relative to the emission seen in G235M, but it is also true that the G235M emission is not symmetrically distributed and its peak corresponds to the maximum emission in G235H, within half a (resampled) pixel. It is possible that there is a residual alignment issue, but it is also possible that the two maps do not match exactly because they are probing slightly different velocity ranges, as discussed below.
@@
The HeII flux reported in Maiolino et al. (2024b), was integrated on a larger velocity range (simply because of the lower resolution and larger spectral channels), and also in a larger aperture (), centred more towards the South relative to the smaller () aperture used in this paper. We therefore re-extract a spectrum on an aperture and velocity range matching the extraction of Maiolino et al. (2024b) – the spectrum is much noisier, but the total flux is , hence matching very well the flux observed in G235M. This indicates that there is some extended HeII flux, diffuse both spatially and in velocity, that is probably not clearly visible in the high-resolution data due to the higher noise affecting the signal distributed on larger velocities. Some of this diffuse emission is actually potentially detected in the HeII maps of Figs.2 and 4, and was potentially seen also in the medium resolution map of Maiolino et al. (2024b). The origin of such diffuse emission, probably with high velocity (of the order of a few hundreds km/s), is not clear. However, we note that Hamel-Bravo et al. (2025) detected broad (655 km/s) HeII emission (without corresponding broad Balmer emission) in the halo of the ultra-metal poor local galaxy SBS 0335-052E; the nature of such broad HeII emission is not clear, but could result from near-pristine gas shocked to high temperature by the outflow driven by the central galaxy. A similar scenario could be happening in the vicinity of Hebe: the vigorous star-forming activity may have driven an outflow that is shock-heating the gas to high temperatures. However, the inferred velocities are quite modest ( km/s), so it is unlikely that shocks are relevant (Lecroq et al., 2024). Additionally, we warn that the presence of high-velocity gas is inferred only indirectly from the total flux in the large aperture, and when compared with the medium-resolution data, deeper data would be needed to directly confirm it.

### 24 Mar 2026 v2

@@ Abstract
The detection and characterization of the first population of stars, formed out of pristine clouds of gas, the so-called Population III (PopIII) stars, is one of the most important goals of modern astrophysics. These stars are theorised to form in galactic embryos within the first few hundred million years after the Big Bang (e.g., Abel et al., 2002; Yoshida et al., 2003; Klessen and Glover, 2023). There have been detections of galaxies with extremely low metallicities, in the range (e.g., Vanzella et al., 2023; Fujimoto et al., 2025; Hsiao et al., 2025; Maiolino et al., 2025; Morishita et al., 2025; Nakajima et al., 2025). These could be very low metallicity PopII galaxies, or possibly self-polluted or hybrid PopIII galaxies (Rusta et al., 2025). However, these metallicities are still well above the expected threshold for PopIII, which are generally identified below a critical metallicity , which prevents the formation of low-mass stars because of inefficient gas cooling (e.g., Bromm et al., 2002; Schneider et al., 2002; Bromm and Yoshida, 2011).
@@
In this paper, we report deep high-resolution (G235H; ) NIRSpec IFU observations of the same region targeted in the previous study. These new observations confirm the presence of HeII emission in the halo of GN-z11. Further confirmation is obtained through the detection of H at the same location and same redshift, which is reported in a companion paper (Übler et al., 2026).
A second companion paper, Rusta et al. (2026), presents a theoretical interpretation of our findings on this source in the Pop III scenario, also discussing the implications for the still-unknown mass distribution of Pop III stars.
@@ 2 Observations and data processing
Observations were obtained through the JWST programme ID (PI: Maiolino) using the NIRSpec IFU (Jakobsen et al., 2022; Böker et al., 2022, 2023) with the high resolution grating G235H (R2700), covering a field of view of . The observations were performed between May 15 and 17, 2025, split into two visits. Detectors were read out with the NRSIRS2 pattern, with 25 groups and 2 integrations. A total of 39 dithers were executed, with a medium cycling pattern. Unfortunately, during dither position 18 the telescope guiding camera locked onto a hot pixel; hence, this frame had to be discarded in further processing. After removing this frame, the total exposure time was 38.8 hours. The data reduction uses the JWST Science calibration pipeline v1.15.0 with CRDS jwst_1293.pmap. To improve the quality of the datacube, we performed a number of further data processing steps, in addition to the default pipeline. The customised procedures used for residual cosmic rays snowballs flagging, subtraction of pink noise (1/f noise), and removal for failed open MSA shutters signal are described in detail in Perna et al. (2023). Stage 2 was executed using the bad-pixel self-calibration implemented in the standard pipeline, in which all exposures on a given detector are used to identify and flag bad pixels that may have been missed by the reference bad-pixel mask. For the final stage of the pipeline, outlier rejection was performed following D’Eugenio et al. (2024b), with a rejection threshold at the percentile. Before combination, the cube was resampled to a scale of per spaxel.
@@ 3 Hebe: confirmation of a HeII emitter in the halo of GN-z11
The upper panel of Fig.1 shows the spectrum extracted from a aperture centred at East and North of GN-z11 (RA 12:36:25.520, DEC +62:14:31.934). The spectrum shows an emission feature, detected at 6, which is identified as HeII1640 at (centroid of the line), blue-shifted by about 450 km/s relative to GN-z11. Note that, for this paper, the redshift of GN-z11 is consistently remeasured based on its HeII emission in our cube, and found to be ; this is slightly lower than the redshift given by Bunker et al. (2023) using medium resolution spectroscopy - - but it is also slightly higher than the redshift that would be derived from their HeII medium resolution detection, which would give ; these slight inconsistencies are likely due to the lower spectral resolution of the previous observations.
This independently confirms the previous HeII detection at the same location obtained with medium resolution spectroscopy by Maiolino et al. (2024b).
As already mentioned, we name Hebe this HeII emitting source, which is located at 3 kpc from GN-z11.
@@
Fig.2 shows the HeII map by collapsing the continuum-subtracted spectral channels between 1.8995 m and 1.9007 m (the three central channels of the line, i.e. a velocity interval of 187 km/s), which highlights the location of Hebe. The white circle indicates the location of GN-z11. Fig.3 shows the HeII contours again, but overlaid on the NIRCam RGB image, with the Field of View adjusted to show the location of Hebe relative to GN-z11 and a foreground galaxy at z=2.04, which causes some gravitational lensing (see Appendix C).
@@
As discussed in more detail in Section 6 below, the location and flux of the HeII emission are broadly consistent with the previous detection based on the NIRSpec medium resolution grating (Maiolino et al., 2024b), although the comparison with the latter suggests the presence of some diffuse emission, both spatially and in velocity. Fig.2 shows the compactness of the line emission, which is unresolved ( 400 pc in size) at the NIRSpec resolution. We note some diffuse emission extending towards the SW. Even though its surface brightness is not statistically significant (mostly within of the background fluctuations), a similar diffuse emission was seen in the medium resolution map. This will be discussed again in Section 6. There is also another marginally detected emitter 0.8′′ North of GN-z11; this was also marginally detected in medium resolution map, and will be discussed in Section 6.
@@
The HeII emission has a FWHM of about 110 km/s (deconvolved by the spectral resolution)222Following, e.g., D’Eugenio et al., 2025 or Shajib et al., 2025, we take the spectral resolution as the nominal value (Jakobsen et al., 2022), but at 1.9 m this is divided by a factor of 0.7 (determined empirically in those references), which takes into account that, for compact sources, the size of the source can be comparable to the size of the IFU slices., but this apparent width is actually resulting from two marginally resolved components. This is in contrast with the H emission, which is spectrally unresolved (see Übler et al. 2026; note that the two spectra have similar resolution – at 1.9m and at 5m). We have simultaneously fitted HeII and H with two Gaussian components (C1 and C2), as shown by the dashed lines in Fig.1. The two components are separated by km/s, and each of them is consistent with being spectrally unresolved (. The simultaneous fit indicates that H is associated only with C2, while the component C1 of H is undetected.
@@
We have also attempted a simultaneous fit with a single Gaussian component, which results in a significantly worse fit, as shown in Appendix A.
@@
It is difficult to explore whether there is any small spatial offset between the two components C1 and C2, as the maps of individual spectral channels become too noisy for a proper assessment.
@@
The inferred fluxes and luminosities of the HeII line and its two individual components are reported in Table 1. These include an aperture correction factor of 3.03, estimated from the cubes of two stars (ID 1219 –PI: Luetzgendorf – and ID 1222 –PI: Willott–) processed with the same methodology, at the wavelength of 1.9m. The total flux has been estimated by simply summing up the flux in the central 5 spectral pixels of the line. The flux of the individual components is determined by the simultaneous double Gaussian fit.
@@
As discussed in Section 6, there is evidence for a diffuse component of HeII; in the second part of Table 1 we also provide an estimate of Hebe’s HeII flux after subtracting this more diffuse component333This is done by taking the diffuse flux estimated in Sect.6 and deriving the contribution to the extraction aperture..
@@ 4 Equivalent width constraints
Measuring or constraining the equivalent width of the HeII emission is incredibly important, as it can provide key constraints on the nature of the source. This requires a detection of the continuum emission. For extremely low metallicity galaxies, the main difficulty is the continuum detection rather than the nebular line emission. A clear example is the extremely metal poor galaxy LAP-1 at (Vanzella et al., 2023), where multiple nebular emission lines are clearly detected, while the continuum remains undetected despite deep exposures and lensing magnification (Nakajima et al., 2025). The HeII emitter in GN-z11 has a similar issue, with the additional complication that it is in the proximity of the relatively extended foreground galaxy at z=2.028. As discussed in Maiolino et al. (2024b) (see also Tacchella et al., 2023), modelling and removing the emission from the foreground galaxy leaves some residual flux around the location of Hebe. Yet, there is no clear indication of a continuum point source associated with it, suggesting that the flux is some diffuse residual after the foreground galaxy removal. Thanks to the more precise localization of Hebe, we can now confirm the absence of a continuum detection at the 5 upper limit level of 3.1 nJy in F200W. We derive this value from the same aperture adopted for the spectrum extraction, accounting for a correction for a point-like source. Direct photometry (also corrected for the limited aperture size) provides a flux measurement of nJy. The 5 upper limit was corrected for the contribution of the HeII line to the broad-band measurement to get a spectral continuum upper limit. The rest-frame EW(HeII) corresponding to this continuum is 47 Å. We provide a conservative lower limit on the EW of the C1 and C2 components by applying the same continuum upper limit to each of them. This translates into lower limits of 28 Å and 22 Å, for the C1 ad C2 components, respectively. All these measurements are provided in Table 1.
As we discuss extensively in Section 7, the inferred rest-frame EW(HeII) are higher than 20Å, well above the EW expected for a metal-enriched (PopII) stellar population. If the EW(HeII) were Å then we would have expected the detection of a source with a flux higher than 6 nJy in F200W.
@@ 5 Absence of metal emission lines
The G235H/F170LP setup covers a wavelength range where various metal emission lines typically detected in high-redshift galaxies (e.g. D’Eugenio et al., 2024a; Scholtz et al., 2025a) are expected to be seen at the redshift of Hebe. Specifically, the NIV]1483,1486, CIV1548,1551, OIII]1661,1666, and CIII]1906,1908 doublets, as well as the NIII]1749 multiplet, are all within the spectroscopic range of this setup. Considering that in local and high-z galaxies at least some these lines (if not all of them) are stronger than HeII, at least some of these metal lines should be visible also in Hebe. On the contrary, none of these transitions is detected in Hebe. We report conservative upper limits of the flux of these metal lines relative to HeII. The spectral regions around the location of these lines are shown in Appendix A.
@@
The companion paper by Rusta et al. (2026) provides a theoretical interpretation of our observational findings, showing that these non-detections are very constraining of the nature of the ionizing sources.
@@ 6 Consistency with the medium resolution observation and diffuse HeII emission
Secondly, in Fig.4 we compare the R2700 HeII map obtained in this work, with the R1000 HeII map obtained in Maiolino et al. (2024b) (contours). Note that the G235M map shows significant HeII emission associated with GN-z11, not seen in the G235H map because the velocity range integrated in G235M is larger and includes also some HeII emission from GN-z11 (recall that Hebe is blueshifted by km/s relative to GN-z11). We have astrometrically registered the two images via the GN-z11 continuum. Unfortunately, this leaves some uncertainty in the relative positions. Specifically, in addition to the coarseness of the NIRSpec spatial sampling ( native spaxels), there are only two continuum sources that can be used for the alignment: GN-z11 and the foreground galaxy, which is not point-like but fairly extended. The white circle in the figure illustrates the location of GN-z11, and its size illustrates the uncertainty in aligning the two continua. The HeII emission in G235H appears slightly offset towards the NE relative to the emission seen in G235M, although within the maps relative alignment uncertainty. It is also true that the G235M emission is not symmetrically distributed and its peak corresponds to the maximum emission in G235H, within half a (resampled) pixel. It is possible that there is a residual alignment issue, but it is also possible that the two maps do not match exactly because they are probing slightly different velocity ranges, as discussed below.
@@
The HeII flux reported in Maiolino et al. (2024b), was integrated on a larger velocity range (simply because of the lower resolution and larger spectral channels), and also in a larger aperture (), centred more towards the South relative to the smaller () aperture used in this paper. We therefore re-extract a spectrum on an aperture and velocity range matching the extraction of Maiolino et al. (2024b) – the spectrum is much noisier, but the total flux is , hence matching very well the flux observed in G235M. This indicates that there is some extended HeII flux, diffuse both spatially and in velocity, that is probably not clearly visible in the high-resolution data due to the higher noise affecting the signal distributed on larger velocities. Some of this diffuse emission is actually potentially detected in the HeII maps of Figs.2 and 4, and was potentially seen also in the medium resolution map of Maiolino et al. (2024b). The origin of such diffuse emission, probably with high velocity (of the order of a few hundreds km/s), is not clear. However, we note that Hamel-Bravo et al. (2025) detected broad (655 km/s) HeII emission (without corresponding broad Balmer emission) in the halo of the ultra-metal poor local galaxy SBS 0335-052E; the nature of such broad HeII emission is uncertain, but could result from near-pristine gas shocked to high temperature by the outflow driven by the central galaxy. A similar scenario could be happening in the vicinity of Hebe: the vigorous star-forming activity may have driven an outflow that is shock-heating the gas to high temperatures. However, the inferred velocities are quite modest ( km/s), so it is unlikely that shocks are relevant (Lecroq et al., 2024). Additionally, we warn that the presence of high-velocity gas is inferred only indirectly from the total flux in the large aperture, and when compared with the medium-resolution data, deeper data would be needed to directly confirm it.